I need help with logarithm
5.2=-log 2
---
P

Answer by stanbon(75887)

2=-log (2/P)
---------
log(2/P) = -2
2/P = 10^(-2)
2/P = 1/100
P/2 = 100
P = 200
============
Cheers,
Stan H.